OpenStudy (anonymous):

]Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (anonymous):

Parallel lines have the same slope, therefore, the line you are looking for will have a slope of 1. y - y1 = m(x - x1) -- using slope 1 and points(6,-2) y -(-2) = 1(x - 6) y + 2 = x - 6 y = x - 6 - 2 y = x - 8 <-- your equation